OpenStudy (theghostlycreeper):

What is the value of x? 6x−2(x+4)=12

alones (alones):

Start from simplifying both sides \(\ 6x−2(x+4)=12\) \[\ 4x−8=12\] Than add \(8\) to both sides. \[\ 4x−8+8=12+8\] \[\ 4x=20\]

OpenStudy (missmeow):

I ended up with x = 8.

OpenStudy (missmeow):

But then again, I suck at math.

OpenStudy (theghostlycreeper):

who is right

OpenStudy (missmeow):

AloneS

alones (alones):

Your last step would bedividing both sides by 4. \[\frac{ 4x }{ 4 }=\frac{ 20 }{ 4}\] So the final answer would be \[\ x=5\]

OpenStudy (theghostlycreeper):

What is the value of c? 14=2c−6+3c14=2c−6+3c

alones (alones):

Make a new post hun.